The spray characteristics of a highly poisonous liquid fuel, especially particle diameters, are desired to known. It has been suggested that particle diameters dp, spray outlet speed U, liquid viscosity μ , surface tension Y, density ρ and spray injector outlet diameter D are connected. Design a experiment which will be done with with water and specify the necessary conditions for the experiment . Write the data with mathematical expressions, if necessary, with dimensionless numbers.
Answer #1

we have: dp-flu, et, T,S,D). variables are listed SI units of dimensions of above below: Variable unit Dimension (M, L,T). decomparing powers of M,T&L on both sides: . b=0 (n ) * (T) 1+ a-3b+c=0 (L) 4 :. c=-1 :Tis dp Uºpop! ñia de la Let: Xz=ruaphoc:. lt bio =P 11 (mm) (1) -1-90 a = -19-3b +=0 an (L) .. -|-|-3(-1) + (su CE-1 3 T 32 PUD From 0, @ +3 T,= f ( Tz, 13) de f (